OpenStudy (anonymous):

Use l'hopsital RUle to solve for lim x-->Pi / 4 5(1-tanx)secx. I solve for the equation (by plugging in pi/4) and I don't get zero, however, my answer is still wrong, should I still use the hospital rule?

OpenStudy (tkhunny):

\(\tan\left(\dfrac{\pi}{4}\right) = 1\) \(\sec\left(\dfrac{\pi}{4}\right) = \sqrt{2}\) 1) Never "plug in". Sometimes substitution is approprate. 2) This is not an indeterminate form. Please don't use l'Hopital's Rule. 3) Please supply the ENTIRE problem statement.

OpenStudy (anonymous):

That is the entire problem. and I should've calculated in radian... right?

OpenStudy (tkhunny):

You can calculate in Light Years if you like. It's NOT an indeterminate form.

OpenStudy (tkhunny):

That doesn't mean we can't find the limit. We just can't use (or don't need) l'Hopital's Rule.